OpenStudy (anonymous):

What is the common ratio of the geometric sequence whose second and fourth terms are 6 and 54, respectively? 1 2 3 4 @iGreen @nincompoop @anonymous_user

OpenStudy (anonymous):

@LunyMoony

OpenStudy (cwrw238):

4th term = a1* r^3 2nd term = a1* r whre r = common ratio so r^2 = 4th term / 2nd term = 54 / 6

OpenStudy (cwrw238):

r^2 = 9 r = 3

OpenStudy (anonymous):

you there @LunyMoony

OpenStudy (anonymous):

so the answer would be 3?

OpenStudy (lunymoony):

Yes, sry I passed these a while ago in geometry but never really understood them, so i was browsing the internet to learn them better

OpenStudy (lunymoony):

its 2 or 3, cwrw says it is 3 so I believe it is 3

OpenStudy (anonymous):

thanks :) i have a few more questions can you help with? Which sequence is geometric and has 14 as its fifth term and 12 as the common ratio? …, 1, 12, 14, 18, … …, 14, 12, 1, 2,… …, 172, 136,118, 19, … …, 8, 4, 2, 1, …

OpenStudy (cwrw238):

5th term = a1 * r^4 where a1 = first term and r = common ration

OpenStudy (cwrw238):

none of those sequences have a common ratio of 12 - each term of the series is 12 times the previous if r = 12

OpenStudy (cwrw238):

eg 1, 12, 144 ... has r = 12

OpenStudy (cwrw238):

the only sequence in the list which is a GS is the last one with a common ratio of 1/2

OpenStudy (lunymoony):

cwrw answered that question @lexikaylynn